OpenStudy (anonymous):

Televisions and computer screens are usually advertised based on the length of their diagonals. If the height of a computer screen is 11 in. and the width is 14 in., what is the length of the diagonal? (Round to the nearest inch.)

OpenStudy (barrycarter):

Hint: Pythagorean Theorem

OpenStudy (anonymous):

A2+B2=C2

OpenStudy (barrycarter):

Yup, that's it.

OpenStudy (anonymous):

11^2+14^2=C2 121+196=C2

OpenStudy (barrycarter):

Yup, keep going

OpenStudy (anonymous):

after when i do that what do i do next

OpenStudy (barrycarter):

Well, if c^2 = 317, what does c equal?